OpenStudy (anonymous):

Find any points of discontinuity for the rational function. y = ((x +6)(x + 2)(x + 8))/((x + 9)(x + 7)) x = 6, x = 2, x = 8 x = 9 x = 7 x = –9, x = –7 x = –6, x = –2, x = –8

OpenStudy (anonymous):

\[\frac{ (x +6)(x + 2)(x + 8) }{ (x + 9)(x + 7) }\]

OpenStudy (cgreenwade2000):

These are pretty easy to find actually. In a rational function, we already know that the denominator can not equal 0. So to find a point of discontinuity, we just need to look at the denominator and set it equal to 0.

OpenStudy (cgreenwade2000):

In this case (x+9)(x+7) = 0 So we set each one to equal 0 (x+9) = 0 (x+7) = 0 And the answers will tell us the points of discontinuity.

OpenStudy (anonymous):

@cgreenwade2000 so x = -9 and x = -7

OpenStudy (cgreenwade2000):

Correct. And can you see that if x = -9, the denominator is 0 and if x=-7, then the denominator = 0?